OpenStudy (anonymous):

Solve for x : √3x+1 = 5

OpenStudy (lgbasallote):

\[\huge \sqrt{3x+1} = 5\] square both sides \[\huge \implies 3x + 1 =5^2\] \[\huge \implies 3x + 1 = 25\] \[\huge \implies 3x = 25 - 1\] \[\huge \implies 3x = 24\] \[\huge \implies x = \frac{24}{3}\] simplify to get your answer. does that help?